# Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub

### v1

@@ 4.3. RQ3: Development-Related Activities Involved in Shared ChatGPT Conversations
To address RQ3, which focuses on the development-related activities where developers share ChatGPT conversations, we analyzed the distribution of shared ChatGPT links across the five data sources. This result reveals the breadth of developer interactions potentially facilitated by ChatGPT, as illustrated in Figures 7, 8, and 9.
@@
Although there are various ways to classify software development activities, we considered and adapted the classification of software development activities proposed in SWEBOK (Washizaki et al., 2024) and previous studies (Hou et al., 2023; Fan et al., 2023), which are widely recognized in the field of software engineering. Following this, we categorized the relevant activities involving shared developer-ChatGPT conversations, as shown in Figure 7. The development-related activities were then grouped into two primary categories: Development Activities and Supporting Activities.
@@
Development Activities encompass seven conventional phases of the software development life cycle (SDLC), that is, Requirements Analysis, Software Design, Software Development, Software Testing, Software Deployment, Software Maintenance and Evolution, and Software Management.
@@
Supporting Activities comprise two novel AI-assisted activities that facilitate and support the software development process, i.e., Data Handling and Information Seeking.
@@
To facilitate observation and comparison, we use two scales (0~1200 and 0~180) in Figure 8, which shows the statistical distribution of the developer-ChatGPT conversations we collected from GitHub based on the nine activities. Notably, Software Development is the most frequent activity where developers share ChatGPT conversations, accounting for a total of 1,059 (41.57%) interactions. Most of these shared developer-ChatGPT conversations are from Code (775, 30.43%), Commits (108, 4.2%), and Issues (107, 4.2%), underscoring the deep involvement and widespread use of ChatGPT in modifying and improving codebases. Software Maintenance and Evolution is the second largest development-related activity, with a total of 844 (33.14%) interactions. Note that, in this activity, there is a high volume of Commits (579, 22.73%) and a substantial number of Pull Requests (116, 4.55%), illustrating the ongoing need for developers to refine and optimize existing software leveraging ChatGPT. Compared to the above two dominant activities, Information Seeking is less frequent, yet still involving 172 (6.75%) developer-ChatGPT conversations. Moreover, Software Deployment mainly includes deployment problems in Code, followed by moderate developer-ChatGPT conversations in Commits and Issues. Software Design, Software Testing, and Data Handling show relatively balanced developers-ChatGPT conversations in Code, Commits, Issues, and Pull Requests. Software Management involves ChatGPT links mainly in Commits with Requirements Analysis having the least developer-ChatGPT conversations.
@@
Figure 9 offers a different perspective on a visual distribution heat map, illustrating the development-related activities involving shared ChatGPT conversations across the five sources. Each data source is illustrated with its respective percentage allocation, offering insights into how ChatGPT is utilized across different sources on GitHub.
@@
Code (1,105): The dominant activity where ChatGPT conversations are shared is Software Development, accounting for 70.14% of all activities, while the proportions of Software Maintenance and Evolution (6.06%) and Information Seeking (9.68%) are much lower. This implies that developers most frequently rely on ChatGPT to directly support core development tasks, such as generating and understanding code. Other activities, such as Software Deployment (4.62%), Software Design (3.35%), and Software Testing (3.44%), represent comparatively lower percentages, further reinforcing that most activities in Code on GitHub involve direct code generation rather than less prominent activities like design or testing.
@@
Commits (823): This source reveals a strong focus on Software Maintenance and Evolution (70.35%), followed by Software Development (13.00%), emphasizing ChatGPT’s role in iterative improvement and code refinement through code modifications. This aligns with our intuition, given that GitHub Commits generally indicates that developers are actively working on improving and maintaining the software, committing updates regularly to ensure software stability and improvements. Other activities like Software Management (6.80%) and Software Deployment (4.25%) contribute less, indicating relatively limited engagement with ChatGPT for software management or deployment.
@@
Issues (295): Software Development again ranks highest at 36.61%, followed by Software Maintenance and Evolution (25.08%), indicating ChatGPT’s utility in addressing core development needs like debugging and code refinement. Besides, Information Seeking (9.49%) is a notable activity, highlighting ChatGPT’s role as a tool for knowledge retrieval within collaborative problem-solving. Software Deployment (9.15%) and Software Design (7.80%) show lower percentages, suggesting that developers also turn to ChatGPT for deployment-related troubleshooting and refining design-level decisions during issue resolution.
@@
Pull Requests (266): This source is primarily related to Software Maintenance and Evolution (43.61%), followed by Software Development (18.42%) and Data Handling (12.41%). This distribution indicates that developers often turn to ChatGPT for quality assurance, feature enhancements, and data processing. Other activities show less involvement in pull requests, further highlighting ChatGPT’s role as a supporting tool for ensuring software quality and consistency during code review.
@@
Discussions (58): Software Development again emerges as the leading activity (34.48%), and followed by Information Seeking ranks second at 24.14%, illustrating ChatGPT’s value as an on-demand knowledge source. Furthermore, Software Deployment (15.52%) and Software Maintenance and Evolution (13.79%) follow, where ChatGPT helps to troubleshoot and guide deployment. Software Design (6.90%) also appears occasionally, showing that discussions often involve high-level architecture and planning with ChatGPT’s assistance.
@@ 4.4. RQ4: Tasks of ChatGPT Conversations Involved in Software Development
To answer RQ4, we investigated the specific tasks within developer-ChatGPT interactions during development. Figure 10 presents a Sankey diagram that visually breaks down nine development-related activities (see Section 4.3) into specific SE tasks with fine granularity. The Sankey diagram effectively illustrates the relationships between the five data sources, nine development-related activities, and 39 detailed SE tasks derived from ChatGPT interactions observed on our collected GitHub data, reflecting how developers utilize ChatGPT in their daily workflows.
@@
The most prominent flow originates from Code (1,105) and Software Development (1,059), channeling into Code Generation and Completion (888), which reflects the frequent practice of using ChatGPT’s capability to automate code generation during software development. Similarly, Commits (823) strongly correlate with Software Maintenance and Evolution (844), contributing to Code Modification and Optimization (686) and Code Review and Debugging (184), further reinforcing the focus on iterative code refinement. SE Tasks like Test Case Generation (59) derive from Software Testing (83), suggesting that ChatGPT aids in quality assurance and defect resolution.
@@ 5.1.1. Characteristics of ChatGPT Usage on GitHub
According to the results of RQ1, the pronounced dominance of Code in shared ChatGPT interactions (43.4%) indicates ChatGPT’s prevalence in code-related tasks like generating code snippets and providing code explanations. Similarly, Commits (32.3%), holding the second largest proportion, reflects developers’ trust in ChatGPT to refine and validate code changes before integration. The sharp increase in ChatGPT’s usage following the release of its sharing feature (peaking around August 2023, just three months later after the release) illustrates the rapid adaptability of developers in embracing AI tools that can be smoothly integrated into their existing workflows. In contrast, the relatively limited engagement in Discussions suggests that developers tend to favor using ChatGPT for structured, task-specific interactions over open-ended conversations, aligning with the efficiency-focused demands of modern software development practices (Vaillant et al., 2024). Interestingly, the usage of ChatGPT’s sharing feature peaked in August 2023, after which the growth rate began to decline. This trend likely reflects an initial surge in user engagement driven by the novelty effect following the feature’s launch, which gradually tapered off as the novelty diminished.
Besides, the analysis of prompt turns distributions and descriptions regarding the shared ChatGPT links reveals a versatile integration of ChatGPT into the development process. The short and task-focused prompts (e.g., those prompts with only one or two turns) indicate that developers typically seek immediate and concise support for specific issues (e.g., concept interpretation, code modification), which is crucial in fast-paced coding environments. Moreover, the strong tendency to include contextual descriptions of shared ChatGPT links (especially in Commits, Pull Requests, and Discussions) reflects an underlying emphasis on transparency and clarity within collaborative development environments. This practice not only facilitates peer review and knowledge sharing, but also signifies a cultural shift toward embracing AI-assisted insights as an integral part of SDLC. As developers continue to employ GenAI tools like ChatGPT, these tools are well-positioned to enhance both individual efficiency and collective innovation in SE workflows (Vaillant et al., 2024).
@@ 5.1.3. Development-Related Activities Involving Sharing ChatGPT Conversations
The results of RQ3 reveal ChatGPT’s dual role in supporting both SDLC activities (e.g., Software Design) and emerging AI-augmented activities like Information Seeking and Data Handling. The prominence of Software Development (in Code) and Software Maintenance and Evolution (in Commits) highlights ChatGPT’s value in accelerating the coding process and refining existing functionality. However, ChatGPT’s presence across diverse data sources on GitHub, for example, aiding problem-solving in Issues, quality assurance in Pull Requests, and knowledge retrieval in Discussions, demonstrates ChatGPT’s adaptability to context-specific activities. This suggests that ChatGPT is not merely a code generator but a versatile assistant integrated throughout the SDLC phases to streamline SE workflows.
@@ 5.1.4. SE Tasks Involving ChatGPT’s Engagement in Software Development
The findings for RQ4 indicate that ChatGPT is primarily employed for the automation of coding-related tasks, specifically Code Generation & Completion (888 instances) and Code Modification & Optimization (686 instances), underscoring its value in supporting coding and maintenance tasks. In contrast, early-phase tasks like Requirements Elicitation (5 instances) and System Architecture Design (28 instances) exhibit minimal usage of ChatGPT. One possible reason is that developers found ChatGPT unable to fulfill certain tasks owing to its limitations in understanding complex requirements and contributing to system design (Liang et al., 2024; Vaillant et al., 2024; Fan et al., 2023; Hou et al., 2023). Moreover, the distribution regarding the SE tasks in Figure 10 could also confirm such a tendency: developers are predominantly using ChatGPT for on-demand assistance (e.g., coding, maintenance, troubleshooting tasks), rather than as a collaborator for architecture-level planning (e.g., architectural design and optimization). This observation raises a question about whether ChatGPT’s limitations in contextual reasoning (Bang et al., 2023) and domain-specific knowledge constrain its applications in early development stages (Cong-Lem et al., 2025). Consequently, tasks that require intensive human judgment and holistic system understanding (e.g., Requirements Elicitation), remain underexplored in LLM-assisted SE workflows.
@@ 5.2. Implications
For researchers, the findings point to promising opportunities to extend ChatGPT’s utility in certain areas, such as requirements engineering, by tailoring LLMs to domain-specific SE workflows. For example, the concentration of SE tasks in Software Development and Software Maintenance & Evolution implies that ChatGPT is primarily perceived as a coding assistant. Moreover, the findings highlight research opportunities to integrate ChatGPT in certain tasks (e.g., System Architecture Design), thereby improving its capability to boost productivity in underexploited areas. Besides, the study underscores ChatGPT’s transformative potential - not only in reshaping how developers engage in SE tasks, but also in redefining what constitutes a development-related activity in the AI era. This evolution calls for foundational knowledge in software engineering, such as SWEBOK (Washizaki et al., 2024), to adapt alongside emerging GenAI-driven software development practices.
Nevertheless, facing the rapid advances in the field of intelligent software engineering (a.k.a AI4SE) (Hou et al., 2023; Gao et al., 2025), our SE research community should develop new methods, standards, and ethical guidelines that integrate intelligent AI tools or components into the SDLC, to ensure that human-AI collaboration improves code quality, developer productivity, and system maintainability (Gao et al., 2025; Terragni et al., 2024).
@@
Our results reveal that developers utilize ChatGPT in various development-related activities (particularly Software Development and Software Maintenance & Evolution). However, errors in ChatGPT’s outputs could potentially introduce inefficiencies or vulnerabilities in software systems (Liu et al., 2024b, a). Moreover, excessive cognitive reliance on GenAI tools might hinder the development of essential manual problem-solving skills, especially for junior developers. Given the potential of errors and biases in LLM-generated content (Chen et al., 2025; Li et al., 2024), it is imperative to establish robust safeguards, such as human oversight, rigorous validation processes, and strict adherence to ethical and legal standards. These measures are critical to maintaining the quality, security, and trustworthiness of AI-assisted development.

### v2

@@ 4.3. RQ3: Development-Related Activities Involved in Shared ChatGPT Conversations
To address RQ3, which focuses on the development-related activities where developers share ChatGPT conversations, we analyzed the distribution of shared ChatGPT links across the five data sources. Specifically, we applied the Constant Comparison method (Charmaz, 2014; Stol et al., 2016) to establish a categorization of the development-related activities, considering all extracted data items including filenames, path names, and other metadata (see Section 3.5). For example, requirements and user stories are categorized as “Requirement Analysis”; explicit change requests and regression fixes discussions are categorized as “Software Maintenance and Evolution”; and path names containing “xxx/testing” or mentions of unit tests are classified as “Software Testing”. Besides, software documentation includes various documents which can be integrated in lifecycle activities (e.g., requirements specifications, architecture and design documents, test reports). Considering that shared ChatGPT links also serve as a form of documentation recording developer-ChatGPT interactions during software development, we chose not to treat software documentation as a separate activity. Otherwise, all the shared ChatGPT links would fall under this documentation activity. The results of RQ3 reveal the breadth of developer interactions potentially facilitated by ChatGPT, as illustrated in Figures 8, 9, and 10.
@@
Although there are various ways to categorize development-related activities, we considered and adapted the classification proposed in SWEBOK (Washizaki et al., 2024) and previous studies (Hou et al., 2024; Fan et al., 2023), which are widely recognized in the field of software engineering. Following this, we categorized the relevant activities involving shared developer-ChatGPT conversations, as shown in Figure 8. The development-related activities were then grouped into two primary categories: Lifecycle Activities and Supporting Activities.
@@
Lifecycle Activities encompass seven conventional phases of the software development lifecycle (SDLC), that is, Requirements Analysis, Software Design, Software Implementation, Software Testing, Software Deployment, Software Maintenance and Evolution, and Software Management.
@@
Supporting Activities include cross-phase operations (e.g., data handling) that assist and support the SDLC phases and are not tied to a specific SDLC phase. These supporting activities comprise two types of novel AI-assisted activities that facilitate and support the software development process, i.e., Data Handling and Information Seeking.
@@
To facilitate observation and comparison, we use two scales (0~1200 and 0~180) in Figure 9, which shows the distribution of the developer-ChatGPT conversations we collected from GitHub based on the nine activities. Notably, Software Implementation is the most frequent activity where developers share ChatGPT conversations, accounting for a total of 1,059 (41.57%) interactions. Most of these shared developer-ChatGPT conversations are from Code (775, 30.43%), Commits (108, 4.2%), and Issues (107, 4.2%), underscoring the deep involvement and widespread use of ChatGPT in modifying and improving codebases. Software Maintenance and Evolution is the second largest development-related activity, with a total of 844 (33.14%) interactions. Note that, in this activity, there is a high volume of Commits (579, 22.73%) and a substantial number of Pull Requests (116, 4.55%), illustrating the ongoing need for developers to refine and optimize existing software leveraging ChatGPT. Compared to the above two dominant activities, Information Seeking is less frequent, yet still involving 172 (6.75%) developer-ChatGPT conversations. Moreover, Software Deployment mainly includes deployment problems in Code, followed by moderate developer-ChatGPT conversations in Commits and Issues. Software Design, Software Testing, and Data Handling show relatively balanced developers-ChatGPT conversations in Code, Commits, Issues, and Pull Requests. Software Management involves ChatGPT links mainly in Commits with Requirements Analysis having the least developer-ChatGPT conversations.
@@
To provide a different and more in-depth perspective on the development-related activities involving shared ChatGPT conversations across the five sources on GitHub, Figure 10 offers this perspective on a visual distribution heat map. This heat map depicts the percentage allocation for each data source, thereby providing detailed insights into how ChatGPT links are shared across the five sources.
@@
Code (1,105): The dominant activity where ChatGPT conversations are shared is Software Implementation, accounting for 70.14% of all activities, while the proportions of Software Maintenance and Evolution (6.06%) and Information Seeking (9.68%) are much lower. This implies that developers most frequently rely on ChatGPT to directly support core development tasks, such as generating and understanding code. Other activities, such as Software Deployment (4.62%), Software Design (3.35%), and Software Testing (3.44%), represent comparatively lower percentages, further reinforcing that most activities in Code on GitHub involve direct code generation rather than less prominent activities like design or testing.
@@
Commits (823): This source reveals a strong focus on Software Maintenance and Evolution (70.35%), followed by Software Implementation (13.00%), emphasizing ChatGPT’s role in iterative improvement and code refinement through code modifications. This aligns with our intuition, given that GitHub Commits generally indicates that developers are actively working on improving and maintaining the software, committing updates regularly to ensure software stability and improvements. Other activities like Software Management (6.80%) and Software Deployment (4.25%) contribute less, indicating relatively limited engagement with ChatGPT for software management or deployment.
@@
Issues (295): Software Implementation again ranks highest at 36.61%, followed by Software Maintenance and Evolution (25.08%), indicating ChatGPT’s utility in addressing core development needs like debugging and code refinement. Besides, Information Seeking (9.49%) is a notable activity, highlighting ChatGPT’s role as a tool for knowledge retrieval within collaborative problem-solving. Software Deployment (9.15%) and Software Design (7.80%) show lower percentages, suggesting that developers also turn to ChatGPT for deployment-related troubleshooting and refining design-level decisions during issue resolution.
@@
Pull Requests (266): This source is primarily related to Software Maintenance and Evolution (43.61%), followed by Software Implementation (18.42%) and Data Handling (12.41%). This distribution indicates that developers often turn to ChatGPT for quality assurance, feature enhancements, and data processing. Other activities show less involvement in pull requests, further highlighting ChatGPT’s role as a supporting tool for ensuring software quality and consistency during code review.
@@
Discussions (58): Software Implementation again emerges as the leading activity (34.48%), and followed by Information Seeking ranks second at 24.14%, illustrating ChatGPT’s value as an on-demand knowledge source. Furthermore, Software Deployment (15.52%) and Software Maintenance and Evolution (13.79%) follow, where ChatGPT helps to troubleshoot and guide deployment. Software Design (6.90%) also appears occasionally, showing that discussions often involve high-level architecture and planning with ChatGPT’s assistance.
@@ 4.4. RQ4: Tasks of ChatGPT Conversations Involved in Software Development
To answer RQ4, we investigated the specific tasks within developer-ChatGPT interactions during development. Here, we adopted a broader view of tasks by conducting an in-depth analysis of developers’ intents, which encompass both action-oriented tasks (e.g., code generation) and knowledge-oriented goals (e.g., understanding specialized tools and techniques). While the latter may not constitute a task in the strict operational sense, we considered them equally relevant for characterizing developers’ practical use of ChatGPT, since seeking such knowledge is often an essential step toward completing a development task. Figure 11 presents a Sankey diagram that visually decomposes nine development-related activities (see Section 4.3) into fine-grained SE tasks. This visualization highlights the relationships between the five data sources, nine development-related activities, and 39 detailed SE tasks derived from ChatGPT interactions in our collected GitHub data, thereby reflecting how developers utilize ChatGPT in their daily workflows.
@@
The most prominent flow originates from Code (1,105) and Software Implementation (1,059), channeling into Code Generation and Completion (888), which reflects the frequent practice of using ChatGPT’s capability to automate code generation during software development. Similarly, Commits (823) strongly correlate with Software Maintenance and Evolution (844), contributing to Code Modification and Optimization (686) and Code Review and Debugging (184), further reinforcing the focus on iterative code refinement. SE Tasks like Test Case Generation (59) derive from Software Testing (83), suggesting that ChatGPT aids in quality assurance and defect resolution.
@@ 5.1.1. Characteristics of ChatGPT Usage on GitHub
According to the results of RQ1, the dominance of Code (43.4%, 1,105) and Commits (32.3%, 823) in shared ChatGPT interactions indicates that developers might predominantly employ ChatGPT as a technical assistant embedded in the core development tasks (e.g., writing, revising, debugging, and integrating code). The results also reflect that ChatGPT’s role in knowledge-sharing activities such as Issue discussions (11.6%, 295) remains limited. The sharp increase in ChatGPT usage following the release of its sharing feature (peaking around August 2023, just three months later after the release) illustrates the rapid adaptability of developers in embracing AI tools that can be smoothly integrated into their existing workflows. In contrast, the relatively limited engagement in Discussions suggests that developers tend to favor using ChatGPT for structured, task-specific interactions over open-ended conversations, aligning with the efficiency-focused demands of modern software development practices (Vaillant et al., 2024). Interestingly, the usage of ChatGPT’s sharing feature peaked in August 2023, after which the growth rate began to decline. This trend likely reflects an initial surge in user engagement driven by the novelty effect following the feature’s launch, which gradually tapered off as the novelty diminished.
Besides, the analysis of prompt turns distributions and descriptions regarding the shared ChatGPT links reveals a versatile integration of ChatGPT into the development process. The short and task-focused prompts (e.g., those prompts with only one or two turns) indicate that developers typically seek immediate and concise suggestions for specific issues (e.g., concept interpretation, code modification), which is crucial in fast-paced coding environments. Moreover, the strong tendency to include contextual descriptions of shared ChatGPT links (especially in Commits, Pull Requests, and Discussions) reflects an underlying emphasis on transparency and clarity within collaborative development environments. This practice not only facilitates peer review and knowledge sharing, but also signifies a cultural shift toward embracing AI-assisted insights as an integral part of SDLC. As developers continue to employ GenAI tools like ChatGPT, these tools are well-positioned to enhance both individual efficiency and collective innovation in SE workflows (Vaillant et al., 2024).
@@ 5.1.3. Development-Related Activities Involving Sharing ChatGPT Conversations
The results of RQ3 reveal ChatGPT’s dual role in supporting both SDLC activities (e.g., Software Design) and emerging AI-augmented activities like Information Seeking and Data Handling. The prominence of Software Implementation (in Code) and Software Maintenance and Evolution (in Commits) highlights ChatGPT’s value in accelerating the coding process and refining existing functionality. However, ChatGPT’s presence across diverse data sources on GitHub, for example, aiding problem-solving in Issues, quality assurance in Pull Requests, and knowledge retrieval in Discussions, demonstrates ChatGPT’s adaptability to context-specific activities. This suggests that ChatGPT is not merely a code generator but a versatile assistant integrated throughout the SDLC phases to streamline SE workflows.
@@ 5.1.4. SE Tasks Involving ChatGPT’s Engagement in Software Development
The findings for RQ4 indicate that ChatGPT is primarily employed for the automation of coding-related tasks, specifically Code Generation & Completion (888 instances) and Code Modification & Optimization (686 instances), underscoring its value in facilitating coding and maintenance tasks. In contrast, early-phase tasks like Requirements Elicitation (5 instances) and System Architecture Design (28 instances) exhibit minimal usage of ChatGPT. One possible reason is that developers found ChatGPT unable to fulfill certain tasks owing to its limitations in understanding complex requirements and contributing to system design (Liang et al., 2024; Vaillant et al., 2024; Fan et al., 2023; Hou et al., 2024). Moreover, the distribution regarding the SE tasks in Figure 11 could also confirm such a tendency: developers are predominantly using ChatGPT for on-demand assistance (e.g., coding, maintenance, troubleshooting tasks), rather than as a collaborator for architecture-level planning (e.g., architectural design and optimization). This observation raises a question about whether ChatGPT’s limitations in contextual reasoning (Bang et al., 2023) and domain-specific knowledge constrain its applications in early development stages (Cong-Lem et al., 2025). Consequently, tasks that require intensive human judgment and holistic system understanding (e.g., Requirements Elicitation), remain underexplored in LLM-assisted SE workflows.
@@ 5.2. Implications
For researchers, the findings point to promising opportunities to extend ChatGPT’s utility in certain areas, such as requirements engineering, by tailoring LLMs to domain-specific SE workflows. For example, the concentration of SE tasks in Software Implementation and Software Maintenance & Evolution implies that ChatGPT is primarily perceived as a coding assistant. Moreover, the findings highlight research opportunities to integrate ChatGPT in certain tasks (e.g., System Architecture Design), thereby improving its capability to boost productivity in underexploited areas. Besides, the study underscores ChatGPT’s transformative potential - not only in reshaping how developers engage in SE tasks, but also in redefining what constitutes a development-related activity in the AI era. This evolution calls for foundational knowledge in software engineering, such as SWEBOK (Washizaki et al., 2024), to adapt alongside emerging GenAI-driven software development practices.
Nevertheless, facing the rapid advances in the field of intelligent software engineering (a.k.a AI4SE) (Hou et al., 2024; Gao et al., 2025), our SE research community should develop new methods, standards, and ethical guidelines that integrate intelligent AI tools or components into the SDLC, to ensure that human-AI collaboration improves code quality, developer productivity, and system maintainability (Gao et al., 2025; Terragni et al., 2024).
@@
Our results reveal that developers utilize ChatGPT in various development-related activities (particularly Software Implementation and Software Maintenance & Evolution). However, errors in ChatGPT’s outputs could potentially introduce inefficiencies or vulnerabilities in software systems (Liu et al., 2024b, a). Moreover, excessive cognitive reliance on GenAI tools might hinder the development of essential manual problem-solving skills, especially for junior developers. Given the potential of errors and biases in LLM-generated content (Chen et al., 2025; Li et al., 2024), it is imperative to establish robust safeguards, such as human oversight, rigorous validation processes, and strict adherence to ethical and legal standards. These measures are critical to maintaining the quality, security, and trustworthiness of AI-assisted development.